User name lkaufman
Log entry time 22:43:36 on June26,2004
Entry number 127033
keyword=Source studes/New PC voltages, IHWP OUT
While the accelerator was down, and Hall A was in a controlled access, we asked if we could have beam in the injector in order to take a PITA scan with the new IHWP setting and the new cathode spot.
Run 2449 - PITA scan, IHWP out, RHWP = 2250, A IA = 5 (after event 2600), only Hall A laser on at 32 uA in injector
Run 2450 - PITA scan, IHWP out, RHWP = 2250, C IA = 0.906 (forgot to set to 5), only Hall C laser on at 30-34 uA in injector
Run 2451 - Current scan, New Pockels cell voltages to zero A_Q (listed below), IHWP out, RHWP = 2250, IAs are as above
Run 2452 - coda failed, restart coda
Run 2453 - coda failed again
kcoda, then restart everything
Run 2454 - RHWP scan with only Hall C beam on in injector, only 3 pts, IHWP out.
MCC was ready to take beam back, and Hall A was back in beam permit; so we ended our tests. We set the RHWP back to 2250 and set the Hall C IA to 5.73.
From Run 2449: PITA slope = -3.24 ppm/V or -194.4 ppm/DAC V.
A_Q offset = 121.46 ppm, so new PC voltages:
POS = 4.290
NEG = 7.503
These new voltages were set before run 2451.
The current scan (run 2451) was a scan of different Hall A and C laser powers on cathode in order to pull off different amounts of current. We would like to study the charge asymmetry (and position differences) we see with these different settings. Perhaps we can get some insight to what cathode effects we might be seeing on our beam.
In closing, the new source settings are:
RHWP = 2250
PC POS = 4.290
PC NEG = 7.503
C IA = 5.73